# Learning the Domain Specific Inverse NUFFT for Accelerated Spiral MRI using Diffusion Models

### v1

@@ 1 Introduction
Considering the latter, acquiring fewer measurments in k-space rewards a decrease in scan time proportional the the ratio of undersampling, but comes at the cost of image quality. Sampling below the Nyquist limit introduces ambiguities during reconstruction which manifest as artifacts in the final image. Reconstruction algorithms must therefore leverage additional information, including multicoil data in the case of parallel imaging, and sparse priors, in the case of compressed sensing, in order to resolve these ambiguities [3, 4].
A third more recent approach to undersampled MRI reconstruction lies in deep learning methods, which essentially learn a set of image priors and use these to regularize solutions to the ill-posed reconstruction problem [5]. Within this category, diffusion models stand out for producing state-of-the art results on image reconstruction tasks for faster scanning, motion correction, noise reduction, and others [6, 7, 8, 9]. Despite this, the vast majority of deep learning approaches, and to our knowledge, all diffusion-based approaches to image reconstruction, focus on Cartesian-sampled MRI, missing out on potential acceleration gains attained by using more efficient non-Cartesian sampling trajectories. This work addresses this gap by introducing a diffusion model-based method for trajectory-agnostic undersampled MRI reconstruction of multicoil spiral MRI. Contributions:

### v2

@@ 1 Introduction
Considering the latter, acquiring fewer measurments in k-space rewards a decrease in scan time proportional to the ratio of undersampling, but comes at the cost of image quality. Sampling below the Nyquist limit introduces ambiguities during reconstruction which manifest as artifacts in the final image. Reconstruction algorithms must therefore leverage additional information, including multicoil data in the case of parallel imaging, and sparse priors, in the case of compressed sensing, in order to resolve these ambiguities [3, 4].
A third more recent approach to undersampled MRI reconstruction lies in deep learning methods, which essentially learn a set of image priors and use these to regularize solutions to the ill-posed reconstruction problem [5]. Within this category, diffusion models stand out for producing state-of-the art results on image reconstruction tasks for faster scanning, motion correction, noise reduction, and others [6, 7, 8, 9]. Despite this, the vast majority of deep learning approaches, and to our knowledge, all diffusion-based approaches to image reconstruction, focus on Cartesian-sampled MRI, missing out on potential acceleration gains attained by using more efficient non-Cartesian sampling trajectories. This work addresses this gap by introducing a diffusion model-based method for trajectory-agnostic undersampled reconstruction of multicoil spiral MRI. Contributions: